Let a,b and C Be three mutually and exhaustive events,find p(b),if 1/3P(c)-1/2P(A)=p(b) please give the answer

Answer by greenestamps(13198)



Not enough information....

Let the probabilities of A, B, and C be x, y, and z, respectively. Then

x+y+z = 1 (the events are exhaustive)

(1/3)z-(1/2)x = y

That is two equations in three unknowns; it is impossible to find a unique solution. There is an infinite number of solutions.